OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation open parentheses x minus 3 close parentheses squared over 4 end fraction minus open parentheses y plus 7 close parentheses squared over 1 = 1?

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@abb0t @zepdrix

OpenStudy (anonymous):

ur vertices is 3 and -7

OpenStudy (anonymous):

the standar form of a hyperbola is (x-h)/a^2 - (y-k)/b^2 = 1.... the coordinates of ur vertices is h and k...

OpenStudy (anonymous):

im sorry what?

OpenStudy (anonymous):

\[\frac{ (x-h)^2 }{ a^2 }-\frac{ (y-k)^2 }{ b^2 } = 1\] it is the standard form of hyperbola that opens right and left.... so when u substitute ur equation in thgis form it becomes \[\frac{ (x-(-3)^2 }{ 4^2 } - \frac{ (y-(7))^2 }{ 1^2} = 1\] so ur h becomes +3 and ur k becomes -7... got it.?